OpenStudy (anonymous):

PLEASE HELP ME!!?? the lateral area of a cone is 612 pi cm^2. the radius is 43 cm. What is the slant height to the nearest tenth of a centimeter? 4.5 7.1 9.1 14.2

OpenStudy (anonymous):

OMG! I love Pi!!!!! 1) Circumferrence of the base= =2(pie)r=74(pie)=74pie lateral area=1/2(slant height)(74pie)=612pie slant height=612 x 2xpie /74pie=16.54 ====================================== 2) total area=lateral area + base area slant height^2=43^2 + (14/2)^2 slant height=43.566 meters base circumference=14pie lateral area=43.566x(1/2)(14pie)=958.066 m^2 base area=7^2xpie=153.93 m^2 total area=1112 m^2

OpenStudy (anonymous):

hmmm.... are you sure? 16.54 isn't even one of the answers

OpenStudy (anonymous):

How does your school teach it dear?

OpenStudy (anonymous):

different, or it seems like that. Idk im just trying to get help, im not seeing this as being the answer from one of the choices. what do you think?

OpenStudy (anonymous):

so you think it might be 14.2 since that's closest to your answer?

OpenStudy (anonymous):

It may. It depends

OpenStudy (anonymous):

well how else am I supposed to find the slant height? you did it wrong or something. it looks good, but that's just not one of the answers. im not sure what you did

OpenStudy (anonymous):

No. Different schools teach in different ways, resulting in different answers. You must display what your solutions are and how they taught it to you in the lesson.

OpenStudy (anonymous):

do you not see the solutions I listed?